OpenStudy (anonymous):

kindly solve this in special products: (1/2r^2-1/4r^2-r^3)^2

OpenStudy (anonymous):

\[(\frac{1}{2r^2}-\frac{1}{4r^2}-r^3)^2\] is this what you want?

OpenStudy (anonymous):

because if so before you square you can combine the first two terms as \[(\frac{1}{2r^2}-\frac{1}{4r^2}-r^3)^2=(\frac{1}{4r^2}-r^3)^2\]

OpenStudy (anonymous):

\[(1/2r^3-1/4r^2-r^3)^2\]

OpenStudy (anonymous):

thats the right given.. hehehe

OpenStudy (anonymous):

oooooooooooooooh

OpenStudy (anonymous):

sorry, .my bad:(

OpenStudy (anonymous):

that's ok but what are you supposed to do with this? multiply it out?

OpenStudy (anonymous):

yes, but using special products..

OpenStudy (anonymous):

i don't know what that means in this case. you have \[(a+b+c)^2=a^2+2 a b+2 a c+b^2+2 b c+c^2\] is this the "special product" you are supposed to use?

OpenStudy (anonymous):

in whatever order you choose. so if you do it here i think the substitution will give \[r^6+\frac{1}{4 r^6}-\frac{1}{4 r^5}+\frac{1}{16 r^4}+\frac{r}{2}-1\]

OpenStudy (anonymous):

how did you get that?